A spring have spring constant k is connected to two masses m and 2m and force of same magnitude are applied on both masses the time period of oscillation of the system is 

Grade:12

1 Answers

Vikas TU
4 years ago
When two blocks are connected to to ends of a spring same spring the formula for time period changes to T = 2π (μ/k)^1/2where, k = spring constantμ = reduced mass = m1m2(m1+m2)in this case μ = 2m*m/(2m+m)μ = 2m^2/3μ = 2m/3so time periodT= 2π (2m/3k)^1/2